### v2

# Measuring Human Contribution in AI-Assisted Content Generation

Abstract
With the growing prevalence of generative artificial intelligence (AI), an increasing amount of content is no longer exclusively generated by humans but by generative AI models with human guidance. This shift presents notable challenges for the delineation of originality due to the varying degrees of human contribution in AI-assisted works. This study raises the research question of measuring human contribution in AI-assisted content generation and introduces a framework to address this question that is grounded in information theory. By calculating mutual information between human input and AI-assisted output relative to self-information of AI-assisted output, we quantify the proportional information contribution of humans in content generation. Our experimental results demonstrate that the proposed measure effectively discriminates between varying degrees of human contribution across multiple creative domains. We hope that this work lays a foundation for measuring human contributions in AI-assisted content generation in the era of generative AI.
{CJK*}
UTF8gbsn
Figure 1: a. Illustration of AI-assisted content generation, where an AI model is prompted with human input and generates output. b. Overview of the proposed method for measuring human contribution, quantified by the ratio of mutual information between human input and AI-assisted output to the total self-information of the AI-assisted output. c. Outcomes of our proposed measure across various poem generation scenarios using Llama-3, involving varying degrees of human contribution (polishing a human poem, generation with the summary, in other words, key ideas, of a human poem, and generation with a poem title). The center line represents the median, the box limits indicate the upper and lower quartiles, the whiskers extend to 1.5x the interquartile range, and the points are outliers. Our measure effectively differentiates varying degrees of human contribution across the scenarios.
Recent advances in large language models (LLMs) [1, 2, 3, 4, 5] have impacted our personal and working lives in significant ways, most notably by changing the process of content generation.[6, 7] Artificial intelligence (AI) “copilots” have emerged as a new and powerful content production tool across a variety of domains [8, 9, 10], such as lyrics creation [11], office work [9, 10], academic writing [8], etc. Consequently, an increasing amount of new content being generated is no longer solely created by humans but is rather the result of AI-assisted creation [12, 13, 14]. In this new creative modality, humans contribute by providing prompts to AI models, resulting in the generation of “AI-assisted output”, as illustrated in Figure 1a.
This development has raised debates about determining the originality and corresponding regulation of content generated with AI assistance [15, 12]. The varying degrees of human contribution in AI-assisted generation complicate the attribution of intellectual contribution to AI-assisted outputs. This issue is particularly pertinent in fields that prioritize originality, such as education [16], academic research [17, 18, 19], and creative work [20]. For example, universities face a dilemma in whether to ban or embrace AI. Administrators and instructors are concerned that students might use AI to create materials for evaluation with varying levels of originality, potentially compromising educational fairness and effectiveness [21, 22]. Similarly, there is a growing debate, underscored by notable incidents [23, 24], concerning the copyright eligibility of AI-assisted works [20, 25, 26].
At the two extreme ends of the human–AI contribution spectrum, the attribution of originality is relatively clear. If a human author simply uses AI to polish their document, it should be considered the result of the author’s own work. Conversely, if a human uses a short, less-informative prompt to generate a large amount of text, it will not reflect much of the human’s intellectual conception. However, there remains a substantial grey area between these two extremes, in which determining originality requires insight into the degree of human contribution during the AI-assisted generation process. Hence, there is an urgent need for a credible measure by which to evaluate human contribution in AI-assisted content generation.
In this paper, we address the quantification of human contribution in AI-assisted content generation. We begin with the recognition that a major obstacle is the lack of a well-defined perspective, or medium, by which to ascertain the extent to which content output can be attributed to humans rather than the AI tools they have used. Towards this goal, we introduce a new general framework within which we provide a preliminary attempt to quantify human contribution in AI-assisted content generation. Our framework hinges on the concept of information content as a modeling medium. Utilizing principles from information theory [27], as depicted in Figure 1b, our approach centers on the quantification of the proportion of the information content in the AI-assisted output that can be attributed to human input. Specifically, it is a ratio of two quantities. The denominator is the total/unconditional information content (surprisal) in AI-assisted output, calculated as the negative logarithm of the probability of generating the AI-assisted generated content, which we refer to as self-information, $I(\bm{y})$. The numerator, $I(\bm{x};\bm{y})$, is the portion of self-information $I(\bm{y})$ that is shared with the total information content from human input, $I(\bm{x})$, which we define as mutual information. The difference between the two is the conditional self-information in AI-assisted output given user input, $I(\bm{y}|\bm{x})$, calculated as the negative logarithm of the probability of generating the AI-assisted output conditional on human input.
We systematically validate the proposed method as a reliable measure of human contribution by evaluating its effectiveness, domain adaptivity, and model adaptivity. To achieve this, we construct a comprehensive dataset of AI-assisted content generation, encompassing various levels of human contribution, multiple creative domains, and outputs from different LLMs. For instance, Figure 1c illustrates the distribution of the outcomes of our proposed measure for AI-assisted poem generation, across three varying levels of human contribution, ranging from high to low, using the LLM Llama-3. Our proposed measure effectively discriminates between varying degrees of contribution, generally producing lower values for content with less human contribution. Additionally, we investigate the impact of content length, resilience to adaptive attacks, and generalization of our method in evaluation. We further apply our measure to real-world human–AI co-creation data, demonstrating its practical applicability. To broaden the applicability of our human contribution evaluation framework to real-world scenarios—where humans may engage in complex content generation processes or where human input is unknown or impossible to measure—we propose an extension that estimates human contribution based on any text alone, without requiring access to its generation process. In brief, this paper poses a novel research question on quantitatively evaluating human contribution in AI-assisted generation and presents simple yet effective information-based measures as potential solutions.
Results
Table 1: Detailed statistics of the constructed dataset.
  Type Corpus # Content Words # Summary Words # Title Words # Subject Words
Paper Abstract Arxiv 134.24±63.07 68.38±20.59 9.63±3.79 2.87±1.37
News News Articles 532.85±86.26 78.36±16.76 8.91±2.32 4.06±1.00
Patent Abstract HUPD 171.65±24.22 59.89±13.85 8.57±5.28 3.91±0.79
Poem Poetry Foundation 208.18±94.08 48.20±11.27 3.65±2.80 -
Dataset Construction
To verify the validity and reliability of the proposed measure of human contribution, we construct a dataset of AI-assisted generation data with known varying levels of human contribution. Note that obtaining a dataset with clear, real-valued labels poses a significant challenge, as there is no established method for quantifying human contribution when both human input and output are known. This gap motivates us to propose our conceptual framework as a meaningful first step toward addressing this complex issue. Therefore, by design, our dataset spans a very large range of human contribution in AI-assisted output with distinct levels that are hardly controversial. For a comprehensive evaluation, we further vary three factors, beyond the level of human contribution: (1) domains, focusing on those where originality protection is crucial; (2) different LLMs; and (3) different random generation runs. Building this dataset primarily involves two steps: raw information collection and processing and AI-assisted content generation.
Raw Information Collection and Processing: First, we collect and process multi-level information in various domains. Specifically, we sample raw data from public datasets across the following domains: academic paper abstracts, news articles, patent abstracts, and poems. We sample 2,000 entries for each domain. For paper abstracts, each raw data entry includes content, title, and subject; for the other three domains, each raw entry includes content and title. Details of the original dataset and sampling process are provided in Supplemental Materials Section 1.1. We further process the data into a uniform structure with decreasing levels of information: content, summary, title, and subject (except poems, because of less-informative titles), with missing parts of the raw data supplemented using GPT-3.5 [2]. The corresponding statistics are presented in Table 1.
AI-Assisted Content Generation: Next, we generate new content using LLMs with varying levels of human input constructed from the earlier process, categorized as follows: polishing, generation with summary, generation with title, generation with subject (where applicable). These inputs use information corresponding to content, summary, title, and subject, respectively. The detailed prompt constructions are shown in Supplementary Materials Section 1.2. These human inputs represent varying levels of human contribution, from high to low, based on the amount of information provided. To support a comprehensive analysis, we apply different LLMs, including the state-of-the-art open-source LLMs Llama-3 [1] and Mixtral [28] and the chatbot GPT-3.5 [2]. We generate $5$ times for a human input with the temperature set as $0.7$ for diverse outputs.
Figure 2: The distribution of the outcomes of the proposed measure for the constructed dataset. The center line represents the median, the box limits indicate the upper and lower quartiles, the whiskers extend to 1.5x the interquartile range, and the points are outliers. Overall, the proposed measure exhibits the expected trend that lower values are obtained for the generated content with less human contribution.
Human Contribution Evaluation
We evaluate the effectiveness of the proposed measure using the constructed dataset. This section focuses on the original scenario where both the AI-assisted output $\bm{y}$ and human input $\bm{x}$ are known, and the AI model $M_{\theta}$’s output probability is available in evaluation. As a result, the generative probability $p_{\bm{\theta}}(\bm{y})$ and the conditional generative probability $p_{\bm{\theta}}(\bm{y}|\bm{x})$ for generating the content can be calculated. In this context, we can directly apply the previously discussed formulas to compute human contribution. Particularly, as shown in Figure 1b, human contribution is calculated as the ratio of the mutual information between human input and AI-assisted output to the total self-information of the AI-assisted output. The detailed calculation is provided in the Methods section. Real-world scenarios corresponding to this situation include originality authentication, where the human author can provide both the AI model’s generative distribution and the original human input for authentication and evaluation, as well as cases where model service providers directly apply the definition during the generation process to calculate a measurement of human contribution.
Overall Trend Analysis: Figure 2 illustrates the human contribution results of two state-of-the-art open-source LLMs, Llama-3 and Mixtral, on the constructed dataset across various domains. From the results, we make the following observations. First, for each combination of model and data domain, varying levels of human contribution yield different distributions for measured human input in the expected direction: the lower the human author’s informational contribution in AI-assisted generation (from polishing, to generation with summary, to generation with title, and finally to generation with subject), the smaller the proposed metric’s value. This indicates that our proposed measure can effectively distinguish different levels of human contribution in AI-assisted generation, providing useful measurements from an informational perspective.
Second, we observe variability in the outcomes of the proposed measure across different data domains for a specific generation mode (e.g., generation with summary). These differences are reasonable because the same generation mode does not necessarily equate to a similar percentage of human contribution across different domains. For example, in our dataset, the News category consists of full articles with detailed and vivid descriptions, while Paper Abstracts are typically concise paragraph. However, the summaries, titles, and subjects used for generation in our dataset for both categories are of comparable information content. Consequently, when generating a news article versus a paper abstract from a summary, title, or subject, the expected human contribution distributions differ due to the inherent difference in the information content of the data. Therefore, in our evaluation, the primary consideration is to verify whether our measure can consistently reflect the overall pattern of varying levels of human contribution for each generation model and creative domain. This consistency would validate the reliability of our proposed measure for distinguishing different levels of human contribution.
Human-annotation Dataset: To further validate the reliability and consistency of our quantitative analysis, we construct a dataset with human annotations. As discussed earlier, obtaining a dataset with clear, real-valued labels remains a significant challenge, as there is no established method for quantifying human contribution when both human input and output are known. Inspired by evaluation platforms such as Chatbot Arena [29], we collect human annotation data by pairwise comparisons. Specifically, we sample 1,000 pairs of input-output data points generated by Llama-3 across all four domains and categories with the condition that the measured human contribution gap between the two points exceeds 0.1 using our method. This threshold is chosen to ensure the meaningfulness of the comparisons and to reduce ambiguity during annotation. Pairs of input-output data points are so annotated as to reflect which one contains a higher degree of human contribution (annotation instructions are detailed in Supplementary Materials section 1.3). Using this annotated dataset, we evaluate how well the relative magnitudes derived from our measurements align with human judgments in these comparisons. The results indicate that 97.4% of the data exhibit consistency between our measurements and human annotations. Moreover, as shown in Supplementary Figure 2, the inconsistent pairs are predominantly concentrated in cases where the measured human contribution gap between two data points is less than 0.2. This further lends support to the robustness and validity of our measurement, demonstrating its consistency across different domains and setups.
Controlled Experiments: In addition to the varying levels of human contribution present in our constructed dataset, we further validate our method by manipulating the total information content of the output while controlling the information content of the input. By doing so, we aim to verify whether our measurement exhibits the expected property that, when the input is controlled, increasing the output information leads to a lower human contribution (as defined by the ratio of human-contributed information relative to the total information in the output). To achieve this, we adjust the output length requirements in the input prompts while keeping the other input information constant. We assume that much longer outputs generally have higher information content. Specifically, we use Llama-3 to generate AI-assisted outputs of varying lengths from titles by specifying the desired output length in the prompt. Figure 3 shows the results for paper and patent abstracts with different lengths, while the results for news and poems are included in Supplementary Materials Section 2.4. The results align with our expectation: as the AI-assisted output contains more information content, the proportion of human informational contribution relative to the total informational content decreases, as reflected by our measurement of human contribution.
Figure 3: The distribution of the outcomes of the proposed measure for academic paper and patent abstracts of different lengths, generated with titles using Llama-3. The center line represents the median, the box limits indicate the upper and lower quartiles, the whiskers extend to 1.5x the interquartile range, and the points are outliers. Overall, the results align with our expectation that with the same human input information the longer the AI-assisted output, the smaller the measured human contribution.
Figure 4: The distribution of the proposed measurement outcomes for academic paper and patent abstracts, generated using different temperature settings with Llama-3. The center line represents the median, the box limits indicate the upper and lower quartiles, the whiskers extend to 1.5x the interquartile range, and the points are outliers. Overall, the results align with our expectation that larger temperature leads to smaller measured human contribution.
Impact of Generative Model Temperature
Temperature refers to the parameter that controls the randomness of output in a generative model. During the generation process, a lower temperature makes the model more deterministic by favoring high-probability tokens, while a higher temperature introduces greater variability, resulting in more diverse but less predictable outputs. In this section, we study the impact of generative model temperature on the outcomes of our measurement. While the “human effort” (human input) remains unchanged across different temperature settings, we expect human contribution—defined as the proportion of informational content in the output attributable to human input—to decrease as the temperature increases. This is because the output becomes less dependent on the input information and increasingly reflects the model’s explorative generation. In extreme cases, when the temperature approaches infinity, the output of the language model would become entirely independent of the human input, resulting in an expected human contribution of zero.
Figure 4 demonstrates the comparison between temperature settings from 0.3 to 0.9 during the generation with Llama-3 for paper and patent abstracts, while the results for news and poems are included in Supplementary Materials Section 2.5. We can observe that generally, a higher temperature leads to a smaller measured human contribution across different generation categories. This validates that our measurement accurately reflects the diminishing influence of human input to AI-assisted output with increasing temperature of the generative model.
Figure 5: The distribution of the outcomes of the proposed measure for the constructed dataset of news with and without adaptive attacks using Llama-3. The center line represents the median, the box limits indicate the upper and lower quartiles, the whiskers extend to 1.5x the interquartile range, and the points are outliers. Overall, the adaptive attacks have little to no influence on the measurement outcomes.
Resilience to Adaptive Attacks
We further investigate whether adaptive attacks could be employed in real-world applications to artificially inflate measured human contribution. To this end, we design two adaptive attacks: we separately append two instructions to the original input that do not provide additional information but do guide the AI’s generation process to potentially increase the measured human contribution. These instructions are to 1) always choose words you rarely use and 2) mimic human writing. The first instruction influences the model’s generation probabilities to produce less frequently used words, thereby attempting to increase the perceived information content (surprisal). The second instruction guides the model to generate text that closely resembles human writing, thereby attempting to increase the perceived human contribution.
Figure 5 shows the results of our measure with and without attacks using Llama-3 in the news domain. The results for other domains are included in Supplementary Materials Section 2.6. We can observe that our measure remains robust against the adaptive attacks. This aligns with our expectation, as we measure human contribution by utilizing the ratio of the mutual information between human input and AI-assisted output to self-information of the AI-assisted output itself. These non-informational instructions for manipulating the output do not significantly affect our measure.
Figure 6: The distribution of the outcomes of the proposed measure for the constructed dataset of news articles for various generation models (columns) and surrogate models for measurement (rows). The center line represents the median, the box limits indicate the upper and lower quartiles, the whiskers extend to 1.5x the interquartile range, and the points are outliers. Overall, for each model pair, the proposed measure exhibits the expected pattern that lower measured values are obtained for the generated content with less human contribution.
Generalization of Our Method
In real-world applications, the AI model’s generative probability $p_{\bm{\theta}}$ may not be available. For instance, generative applications like ChatGPT do release generative probabilities to users. This section demonstrates whether a surrogate model with generative probability $p_{\bm{\theta}}^{\prime}$ can be employed for our assessment when the AI model’s generative probability $p_{\bm{\theta}}$ is unknown. Specifically, in this experiment, we use Llama-3 and Mistral as the surrogate models and use their generative probability $p_{\bm{\theta}}^{\prime}$ to assess the content generated by various LLMs (Llama-3, Mistral, and ChatGPT) in the constructed dataset. Note that generation models can include chatbots or models that do not release their generative probabilities, while surrogate models require access to generative probability distributions for evaluation.
Figure 6 illustrates the effectiveness of our approach in the news domain for various combinations of surrogate model (rows) and generation model (columns). Results for other domains are presented in Supplementary Materials Section 2.7. We observe that even without using the original AI model for evaluation, our proposed measure captures the expected trend in human contribution across various surrogate and generative model combinations. This validates the applicability of our measure when generation model information is unavailable. This effectiveness may be attributed to the similarity in generative distributions across LLMs, stemming from the universal knowledge they share during training. The gradient in human contribution across varying levels of human input is far more pronounced than the differences between the distributions themselves, indicating that our method is a robust assessment tool.
Figure 7: a. The distribution of the outcomes of the proposed measure across various classes of real-world AI-assisted generation, sampled from WildChat. The center line represents the median, the box limits indicate the upper and lower quartiles, the whiskers extend to 1.5x the interquartile range, and the points are outliers. The “editing or rewriting” class typically exhibits a higher degree of measured human contribution. b. Case studies on data samples from two distinct classes in WildChat.
Applications to Real-World AI-Assisted Generation
The aforementioned experiments were conducted on a synthetic dataset with known varying levels of human contribution, allowing us to verify the reliability of our measurement method. To test its real-world applicability, we apply our method to real-world scenarios involving user interactions with LLMs. Specifically, we sample cases from the WildChat-1M dataset [30] and classify them using a prompt classification tool [31]. We then sample data from two prompt classes related to AI-assisted generation: “assisting or creative writing” (2,000 entries) and “editing or rewriting” (500 entries), according to their counts in the dataset. The evaluation surrogate model is Llama-3, while the contents were generated with ChatGPT.
Figure 7a demonstrates the overall distribution of measured human contributions across the two classes. We expect that the “editing or rewriting” class will involve more human contribution than “assisting or creative writing.” Consistent with this, the measured human contributions are generally higher for “editing or rewriting.” We present two specific cases for two classes respectively in Figure 7b, with additional cases detailed in Supplementary Materials Section 2.8. Overall, the measured human contributions align with our expectations. For instance, the “editing or rewriting” case is measured as having 92.86% human contribution, while the “assisting or creative writing” case is measured at 16.14%. These distribution and case study results further support the validity of our method in measuring human contribution in real-world AI-assisted generation contexts.
Figure 8: The distribution of the outcomes of the proposed estimation for the constructed dataset. The center line represents the median, the box limits indicate the upper and lower quartiles, the whiskers extend to 1.5x the interquartile range, and the points are outliers. Overall, the proposed estimation exhibits the expected trend that lower values are obtained for the generated content with less human contribution.
Human Contribution Estimation Without Human Input
To better address real-world scenarios where the AI-assisted generation process is unknown, we extend our method to estimate the human contribution of any given text alone without knowledge of human input and output generation processes. Clearly, we have less information now than the situation considered before. This loss of information leads to a degree of uncertainty, as the same output can be generated from inputs with varying levels of contribution. However, we address this problem of uncertainty by redefining our goal to estimate the proportion of information content in the AI-assisted output that is necessarily provided by humans. This is also referred to as the minimum human contribution for human input $\bm{x}$ to plausibly generate the AI-assisted output $\bm{y}$. This focus on minimal human contribution stems from several practical considerations, such as evaluating the unique original informativeness value that LLMs cannot easily generate, as detailed in the Methods section.
To compute this, we adapt the original formula $\frac{I(\bm{x};\bm{y})}{I(\bm{y})}$. While the denominator $I(\bm{y})$ can still be directly calculated, the numerator $I(\bm{x};\bm{y})=I(\bm{y})-I(\bm{y}|\bm{x})$ is estimated as its minimum possible value for a valid hypothetical $\bm{x}$ that plausibly generates $\bm{y}$. Intuitively, if a hypothetical $\bm{x}$ plausibly generates $\bm{y}$, the likelihood of $\bm{y}$ being generated based on $\bm{x}$, $p_{\theta}(\bm{y}\mid\bm{x})$, should be sufficiently high (have a lower bound). This implies that the conditional information (or surprisal) $I(\bm{y}|\bm{x})=-log(p_{\theta}(\bm{y}\mid\bm{x}))$ of generating $\bm{y}$ given $\bm{x}$ should have an upper bound. We leverage this upper bound to compute the minimum possible value of $I(\bm{x};\bm{y})$. The detailed motivation for estimating minimum human contribution, along with the calculation method, is provided in the Methods section.
Overall Trend Analysis: In Figure 8, we present the human contribution estimation results for two state-of-the-art open-source LLMs, Llama-3 and Mixtral, applied to the constructed dataset across various domains. We make the following observations: First, the proposed estimation method produces results that follow the expected trend in our dataset: as the human author’s informational contribution decreases during AI-assisted generation, the estimated metric value correspondingly decreases. Second, as we suppress the diverse real values of $I(\bm{y}\mid\bm{x})$ to an estimated boundary value that is empirically chosen, the resulting distributions do not fully match human contribution evaluations where $\bm{x}$ is known (Figure 2). Finally, since this is an estimation process, there are cases where values fall below zero. These can be interpreted as instances where the corresponding output might plausibly be generated with minimal or negligible human contribution under the minimal human contribution assumption. Overall, our method provides a meaningful solution for estimating human contribution in situations where direct input information and generation processes are unavailable.
Figure 9: a. Illustration of the tested multi-round refinement scenario. The human contribution to the output is expected to decrease from round 1 to round 3, as the output progressively incorporates more information with minimal additional input information from the human. b. The evaluated human contribution relationship across outputs from different rounds. Our evaluation produces results that align with our expectations: the human contributions evaluated from the outputs of the later rounds are consistent with those of the corresponding previous rounds, showing a gradual decrease from round to round.
Multi-Round Generation Analysis: In addition, we further investigate the validity of our evaluation in the context of multi-round refinement. We design an experiment, as illustrated in Figure 9a, where we sample the generation of 1,000 news articles from the original dataset for the first round of generation (covering all generation modes). Subsequently, we add second and third rounds of refinement, requiring the model to continuously incorporate additional information. Intuitively, the proportion of human contribution decreases from the round 1 output to the round 3 output, as the output progressively incorporates more information while minimal additional input from the human is given.
The experimental results are presented in Figure 9b. The left panel shows the relationship between the human contribution evaluated for the round 1 output (x-axis) and the round 2 output (y-axis), while the right panel depicts the relationship between the round 2 output (x-axis) and the round 3 output (y-axis). These figures illustrate changes in human contribution across different rounds of output. The results align with our expectations: there is consistency across rounds, meaning the human contribution differences introduced by various generation modes (e.g., polish, generation with summary, title, or subject) in the first round persist in the subsequent rounds. At the same time, most data points lie below the line $y=x$, indicating that human contribution generally decreases in the later rounds. Statistically, 87.12% of cases show a reduction in human contribution from round 1 to round 2 , and 94.37% show a reduction from round 2 to round 3.
Discussion
Technological advancements in generative AI have significantly altered the content production process, resulting in the generation of a vast amount of AI-assisted content  [11, 10, 8]. This proliferation of AI-assisted generated content poses challenges to the delineation of originality of these works, sparking intense debates regarding the application of AI in areas that prioritize originality, such as education, academic research, and copyright [23, 24, 25, 32, 33, 34]. It has been increasingly recognized that defining the intellectual originality of AI-assisted generated content cannot be approached with a one-size-fits-all solution; instead, it requires consideration of the extent of human contribution to the work [20, 22, 33]. Consequently, to facilitate the further development of generative AI, decision makers will need credible methods to measure human contribution in AI-assisted generation across various contexts.
This study frames this challenge and presents a method to measure human contribution to AI-assisted generation that is grounded in information theory. We propose using information content to quantify the percentage of information in the AI-assisted output that is attributable to human input. By measuring the ratio of mutual information between human input and AI-assisted output to the self-information of AI-assisted output, we quantify the human contribution in AI-assisted generation. Our measure is validated through experiments conducted on multi-domain AI-assisted generation datasets using multiple LLMs. In raising this research question and proposing a new framework, we seek to measure human contribution quantitatively, inspire further research, and help advance the refinement of relevant originality delineation and content regulation in the future.
While it serves to frame the question and provide a preliminary method for measuring human contribution, our work has several limitations. First, the current framework focuses on textual output from LLMs. However, originality issues related to AI-assisted generation are not limited to text; they also extend to images, audio, video, computer code, etc. Incorporating non-textual output raises even more complex problems due to the change in modality between human input and AI-assisted output. We aim to explore originality issues in such scenarios in future research. In other modalities, since data may no longer be discrete, this framework may not be directly applicable. However, as generative models are fundamentally grounded in certain generative probabilistic distributions, we hope that the information theory-based framework proposed in this work can serve as a foundation and inspiration for addressing originality challenges across diverse modalities. Second, caution is warranted when measuring human contribution in AI-assisted content generation in the copyright domain. Human edition, selection, and compilation of AI-assisted content may provide significant creative input, which could be relevant when assessing authorship for copyright purposes. Finally, we acknowledge that “human contribution” can have different meanings. In this work, we focus specifically on measuring informational contribution, while leaving other interpretations, such as prompt-engineering effectiveness, for future exploration.
Ethical and Societal Impact
The objective of this study is to pose a research question and propose a framework for measuring human contribution in AI-assisted content generation. This question and framework aim to facilitate originality delineation in the era of creation with the assistance of AI. Simultaneously, this work seeks to inspire more research on technical methods that can support the enhancement of relevant regulations in the context of widespread AI utilization in various contexts. A potential risk is that in real-world applications of the proposed framework, there might be targeted adaptive attacks aimed at manipulating the results to artificially elevate the assessed level of human contribution. Although this paper examines two adaptive attacks and verifies the robustness of the proposed measure against them, more sophisticated and advanced attacks may arise in real-world scenarios. We hope to further understand and mitigate such risks in future work.
The authors of this paper introduce a method to technically measure human contribution in AI-assist content generation. The method is widely applicable in different situations. However, the paper does not intend to discuss the complex copyright legal and policy issues related to “originality” or “eligibility,” nor it reflect any of Microsoft’s legal and policy positions on the copyright issues.
Methods
Related Work
The research problem most closely related to evaluating human contribution is the detection of content generated by LLMs[35, 36, 37]. As the performance of LLMs continues to improve, the risk of being unable to distinguish between content generated by LLMs and humans becomes increasingly apparent, with attendant threats in security, fraud prevention[38, 39], and academic integrity [40], among other fields [35, 41, 42]. Consequently, researchers are increasingly directing their efforts towards the detection of LLM-generated content, specifically ascertaining whether a given text is primarily the product of AI. These research efforts entail training detection models [43, 44, 45], employing features for zero-shot detection [46, 37, 47], or incorporating specific watermarks during content generation [36, 48, 49].
While the current body of research predominantly focuses on identifying content substantially generated by AI, thus optimized for binary detection, real-world AI-assisted generation often involves varying degrees of human contribution. In many practical application scenarios, it is not sufficient to merely detect content primarily generated by AI, rather it is crucial to discern the extent of human contribution. Therefore, distinct from the detection of AI-generated content, our emphasis is on reliably quantifying human contribution within AI-assisted generation from an informational perspective.
Defining Human Contribution in AI-Assisted Generation
In contrast to a binary classifier determining whether content is primarily generated by AI, our aim is to derive a quantitative measurement indicating the extent of human contribution in AI-assisted content generation. This is a novel and previously unexplored issue. Our core idea revolves around utilizing information content as a medium for gauging the contributions of humans and AI. Particularly, we define human contribution in AI-assisted generation as the ratio of mutual information between human input and AI-assisted output relative to the total self-information of the AI-assisted output, as illustrated in Figure 1b.
In this section, we first introduce related concepts derived from information theory [27]; we then provide our definition of human contribution. In the following definition, we consider an AI model $M_{\theta}$, its generative distribution $p_{\theta}$, human input $\bm{x}$, and AI-assisted output $\bm{y}$. First, we quantify the information content within the generated output $\bm{y}$ through the concept of self-information. Self-information measures the level of surprisal associated with an event, reflecting the probability of that event occurring. In this context, generating content that is less probable is considered more informative event. We represent the self-information of the generated output $\bm{y}$ as follows:
$I(\bm{y})=-log(p_{\theta}(\bm{y})),$ (1)
where $p_{\theta}(\bm{y})$ is the probability that the content $\bm{y}$ is generated without any condition.
On the other hand, when conditioned on human input $\bm{x}$, the information content within the generated output $\bm{y}$ transforms into conditional self-information. Conditional self-information quantifies the information contained in an event, given the occurrence of another event. Here, we represent the conditional self-information of the generated output $\bm{y}$ given “human input is $\bm{x}$” as follows:
$I(\bm{y}\mid\bm{x})=-log(p_{\theta}(\bm{y}\mid\bm{x})),$ (2)
where $p_{\theta}(\bm{y}\mid\bm{x})$ is the probability that the content $\bm{y}$ is generated conditioned on human input $\bm{x}$. It is noteworthy that both information and conditional information are calculated for the specific event “generated output is $\bm{y}$,” rather than the entire generative distribution—which would be difficult to measure due to the vast space of all possible generated contents. This allows the probabilities $p_{\theta}(\bm{y})$ and $p_{\theta}(\bm{y}\mid\bm{x})$ to be easily computed using the corresponding generative probabilities, without and with human input. Specifically, considering the generative process of the autoregressive LLM $M_{\theta}$, these probabilities can be computed by multiplying the probabilities of each token $y_{i}$ being sampled from the generative distribution, conditioned on the previously generated tokens (and the human input $\bm{x}$, if provided):
$p_{\theta}(\bm{y})=\prod_{i=1}^{N}p_{\theta}(y_{i}\mid\bm{y}_{<i}),$ (3)
$p_{\theta}(\bm{y}\mid\bm{x})=\prod_{i=1}^{N}p_{\theta}(y_{i}\mid\bm{y}_{<i},\bm{x}),$ (4)
where $N$ is the total number of tokens in $\bm{y}$..
Based on these two information concepts, we define (point-wise) mutual information between the generated content $\bm{y}$ and the human input $\bm{x}$ as the reduction in surprisal when human input $\bm{x}$ is known for generating content $\bm{y}$. This indicates how much information content in $\bm{y}$ can be attributed to human input $\bm{x}$. Formally, it is defined as:
$I(\bm{x};\bm{y})=I(\bm{y})-I(\bm{y}\mid\bm{x}).$ (5)
Building upon the aforementioned definition of information within the AI-assisted generation process, we proceed to establishing the definition of human contribution in AI-assisted generation.
Definition 1 (Human contribution in AI-assisted generation).
Given an AI model $M_{\theta}$ and human input $\bm{x}$, where $\bm{y}$ represents the AI-assisted generated content, the human contribution $\phi$ is defined as the ratio of mutual information $I(\bm{x},\bm{y})$ to self-information $I(\bm{y})$: $\phi=\frac{I(\bm{x};\bm{y})}{I(\bm{y})}$
This definition of human contribution pertains to the proportion of the information content within the generated output that can be attributed to human input, relative to the total information content of the generated output. This definition can be directly employed for evaluation in scenarios where human input is available and reliable. For instance, model service providers can apply this definition during the generation process to quantify human contributions and issue certified documentation. Moreover, in evidentiary scenarios for originality authentication, human authors can provide reliable human input (certified with appropriate documentation) for evaluation.
Estimating Human Contribution in AI-Assisted Generation Without Human Input
In many scenarios requiring the evaluation of human contribution, the generation process of AI-assisted output $\bm{y}$ can be complex (e.g., involving iterative back-and-forth refinements) or entirely unavailable. To address this challenge, we extend our framework by introducing the human contribution estimation problem and a corresponding solution that is applicable to any text output, without requiring prior knowledge of the human input $\bm{x}$ or the generation process.
Problem Definition: First, we define a research problem aimed at evaluating human contribution without reliance on $\bm{x}$. It is noteworthy that the same output $\bm{y}$ can result from various human inputs $\bm{x}$, which are associated with differing levels of human contribution. For instance, an AI-generated abstract based solely on a subject (representing minimal human contribution) could also be produced through repetition, where human contribution approaches 100%. To establish a well-defined problem, we frame the research challenge as determining the minimum human contribution required for human input $\bm{x}$ to plausibly generate AI-assisted output $\bm{y}$. Specifically, this involves quantifying the proportion of information content within the AI-assisted output that is necessarily provided by humans.
Our focus on minimal human contribution stems from several practical considerations: First, all outputs can theoretically be generated from a prompt with nearly 100% human contribution, such as prompting LLMs for repetition. Estimating minimal human contribution provides a reference range, emphasizing the lower bound of human-contributed information content required to plausibly generate the output. Second, in the context of widespread LLM adoption, minimal human contribution evaluates the unique original informativeness value provided by humans—content that LLMs cannot easily generate. This distinction is critical in various applications, such as curating high-value data for training next-generation models, assessing creative or original works that reflect human ingenuity and effort, and identifying unique contributions in academic peer reviews.
Definition 2 (Minimal Human Contribution in AI-Assisted Generation).
Given an AI model $M_{\theta}$ and the AI-assisted generated content $\bm{y}$, the minimal human contribution $\hat{\phi}$ is defined as the ratio of the minimal mutual information $\hat{I}(\hat{\bm{x}},\bm{y})$—where $\hat{\bm{x}}$ represents the minimally informative human input required to plausibly generate $\bm{y}$—to the self-information $I(\bm{y})$: $\hat{\phi}=\frac{\hat{I}(\hat{\bm{x}};\bm{y})}{I(\bm{y})}.$
Solution: To solve this problem, we first need to establish a reasonable criterion for determining whether $\bm{x}$ could plausibly generate $\bm{y}$. Considering the generative process of an autoregressive LLM, each token $y_{i}$ is sampled from a generative distribution $p_{\theta}(y_{i}\mid\bm{y_{<i}},\bm{x})$ conditioned on the previous tokens and $\bm{x}$. For $\bm{y}$ to be considered plausibly generated by $\bm{x}$, we require that the geometric mean of the conditional probabilities $p_{\theta}(y_{i}\mid\bm{y}_{<i},\bm{x})$ across all tokens exceeds a predefined threshold $\tau$:
$\left(\prod_{i=1}^{N}p_{\theta}(y_{i}\mid\bm{y}_{<i},\bm{x})\right)^{1/N}>\tau,$ (6)
where $N$ is the number of tokens in $\bm{y}$. This condition ensures that, overall, the likelihood of $\bm{y}$ being sampled from the distribution of model $M_{\theta}$ given input $\bm{x}$ at each step is sufficiently large. It prevents any token in $\bm{y}$ from having an excessively low probability of being sampled, maintaining the credibility of $\bm{x}$ as the source of $\bm{y}$. The threshold $\tau$ can be chosen based on the characteristics of the generative distribution of evaluated LLM as well as the requirements to balance between permissive and strict criteria.
Based on this criterion, we can estimate the minimum human contribution $\hat{\phi}=\frac{\hat{I}(\hat{\bm{x}};\bm{y})}{I(\bm{y})}$. The denominator $I(\bm{y})$ can be directly computed; thus, our primary task is to compute a lower bound for the mutual information $I(\bm{x},\bm{y})=I(\bm{y})-I(\bm{y}\mid\bm{x})$ by deriving an upper bound for the conditional self-information $I(\bm{y}\mid\bm{x})$.
According to our plausibility criterion, we have:
$p_{\theta}(\bm{y}\mid\bm{x})=\prod_{i=1}^{N}p_{\theta}(y_{i}\mid\bm{y}_{<i},\bm{x})>\tau^{N},$ (7)
where $N$ is the total number of tokens in $\bm{y}$, and $\tau$ is the predefined plausibility threshold. From this, the upper bound for the conditional self-information $I(\bm{y}\mid\bm{x})$ can be computed as:
$I(\bm{y}\mid\bm{x})=-\log p_{\theta}(\bm{y}\mid\bm{x})<-N\log\tau.$ (8)
This can be interpreted as follows: when $\bm{x}$ plausibly generates $\bm{y}$, the likelihood of $\bm{y}$ being generated based on $\bm{x}$ is sufficiently high, implying that the conditional information (or surprisal) of generating $\bm{y}$ given $\bm{x}$ should be bounded. From this, we derive the mutual information lower bound:
$I(\bm{x};\bm{y})\geq I(\bm{y})+N\log\tau.$ (9)
Finally, we can compute the estimated minimal human contribution as:
$\hat{\phi}=\frac{\hat{I}(\hat{\bm{x}};\bm{y})}{I(\bm{y})}=\frac{I(\bm{y})+N\log\tau}{I(\bm{y})}.$ (10)
This framework enables the estimation of minimal human contribution $\hat{\phi}$ without requiring knowledge of $\bm{x}$. It can be broadly applied to real-world evaluations of human contribution for any given text, providing a practical solution for scenarios where the AI-assisted generation process is unknown.
Experimental Setups
In this section, we detail the versions and parameters of the models we apply in the experiments are as follows. For evaluation, we use Llama-3 (Llama-3.1-8B-Instruct), Mixtral (Mixtral-8x7B-v0.1). In the estimation experiments, the thresholds $\tau$ are set to $0.65$ and $0.7$, respectively, based on their generative distributions. The generation models further include GPT-3.5 (gpt-3.5-turbo-0125) and GPT-4o (gpt-4o-2024-08-06).
Computational Resources
The experiments in this paper were conducted using NVIDIA A100 80GB GPU on the Ubuntu 22.04.4 LTS platform with an AMD EPYC 7V13 64-Core Processor. Using the proposed metric, measuring one sample in the dataset on a single NVIDIA A100 GPU takes, on average, $0.33$ seconds.
Data Availability
Code Availability
The code applied in the experiments is publicly available at https://github.com/xyq7/Human-Contribution-Measurement.
Supplementary Information accompanies this manuscript in the attached supplementary information file.
Competing Interests: The authors declare no competing interests.